The pH of stomach acid is typically found to be around 1.50 (give or take). HCl is the acid accounting for this acidity, and it's used to help digest protein. Human blood is not nearly as acidic. By approximately what factor of 10 less acidic is human blood?

a) 100,00x less acidic

b) 1,000x

c) 1,000,000x

d) 100x

e) 10,000x

I am really not sure how to approach this problem. If anyone could please assist me :)
Answer #1

Correct option: c

pH of human blood = 7.4

or, - log[H3O+] = 7.4

or, [H3O+] = 10-7.4 M

pH of stomach acid (given) = 1.5

or, - log[H3O+] = 1.5

or, [H3O+] = 10-1.5 M

Hence, the acidity of stomach acid =10-1.5/10-7.4 = 105.9 106 times that of the acidity of human blood.

Therefore, human blood is 1,000,000x less acidic.